OpenStudy (anonymous):

can someone please help me with this: Differentiate 2cos7x*cosx when x is pi/12

OpenStudy (callisto):

~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ Recall: Product rule: \[\frac{d}{dx}[f(x)g(x)] = f'(x)g(x)+g'(x)f(x)\] and \[\frac{d}{dx}cos(ax) = -asin(ax)\] ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ \[\frac{d}{dx}(2cos7x*cosx)\]\[=2(cosx\frac{d}{dx}(cos7x)+cos(7x)\frac{d}{dx}(cosx))\] Can you continue from here?